# Dependence of the estimated electric potential in thunderstorms observed at GRAPES-3 on the hadronic interaction generators used in simulations

### v1

@@ Abstract
A potential difference of 1.3 Giga-Volts (GV) was inferred across a thundercloud using data from the GRAPES-3 muon telescope (G3MT). This was the first-ever estimation of gigavolt potential in thunderstorms, confirming prediction of C.T.R. Wilson almost a century ago. To infer the thundercloud potential required acceleration of muons in atmospheric electric field to be incorporated in the Monte Carlo simulation software CORSIKA. The G3MT records over 4 billion muons daily that are grouped into 169 directions covering 2.3 sr sky. This enabled changes as small as 0.1% in the muon intensity on minute timescale, caused by thunderstorms to be accurately measured. But that requires high statistics simulation of muons in thunderstorm electric fields. The CORSIKA offers a choice of several generators for low- (FLUKA, GHEISHA, and UrQMD) and high-energy (SIBYLL, EPOS-LHC, and QGSJETII) hadronic interactions. Since it is unclear which combination of the low- and high-energy generators provides the correct description of hadronic interactions, all nine combinations of generators were explored, and they yielded thundercloud potentials ranging from 1.3 GV to 1.6 GV for the event recorded on 1 December 2014. The result of SIBYLL-FLUKA combination yielded the lowest thundercloud potential of 1.3 GV was reported. Furthermore, another seven major thunderstorm events recorded between April 2011 and December 2020 were analyzed to measure the dependence of their thundercloud potential on the hadronic interaction generators. It is observed that the low-energy generators produce larger variation (14%) in thundercloud potential than the high-energy generators (8%). This probably reflects the fact that the GeV muons are predominantly produced in low-energy (80 GeV) interactions, which effectively magnifies the differences in the meson production cross-sections among the low-energy generators.
@@ 4 Thunderstorm events recorded by G3MT
In addition, by combining the information from the EFMs on the ground, several important mechanical and electrical properties of this thundercloud were estimated. The electric field on the ground generated by the thundercloud was measured by four widely separated (up to 6 km) EFMs installed around the GRAPES-3 experiment. It was observed that the electric field profiles were very similar but were progressively offset in time almost exactly proportional to their relative distance in the East to West direction. Therefore, the spatial separation between the EFM stations and the corresponding time delays could be used to calculate the speed of the thundercloud of 1 km/min. Similarly, the muon deficit in the FOV of G3MT showed a clear movement, indicative of a movement in the direction of the muon deficit. By using muon images with 2-minute exposures, the angular velocity of the thundercloud was estimated to be 6.2∘/min. Next, by combining the linear (1 km/min.) and angular (6.2∘/min) velocities, a thundercloud height of 11.4 km amsl was obtained. This altitude is comparable to the altitude of a typical cumulonimbus cloud floating nearly at the top of the troposphere. From Figure 3a, it can be observed that the muon deficit extended across the entire range of North-South directions, with an angular diameter of about 74.6∘. Of course, this estimated size is limited by the FOV of the G3MT and the actual size of the thundercloud could be much larger. Therefore, all of the estimates reported here provide a lower limit to the actual size of the thundercloud. The area of the thundercloud was estimated to be 380 km2. By assuming the thundercloud to have a dipolar structure with two charged layers (each 2 km thick) and separated by another 2 km distance, the capacitance of this thundercloud was estimated to be 0.85 F. At a peak potential of 1.3 GV, this capacitor would store a charge of 1100 C and a total energy of 720 GJ. Further details of this event can be found in [17]. The GeV muons detected by the G3MT, being relativistic particles, and combined with the fact that they almost traverse through the entire depth of the atmosphere, make them ideal candidates for thunderstorm studies. This may also explain the lack of success of sounding experiments to measure thundercloud potentials beyond 0.13 GV due to their limited coverage area and live-time [12]. In contrast, G3MT operates throughout the year with a sky coverage of 2.3 sr, which monitors a huge volume of the atmosphere with nearly 100% duty cycle. Although the G3MT has been operational for more than two decades, thunderstorm events could be studied in great detail after the installation of EFMs in April 2011. A collection of 487 significant thunderstorm events recorded from 2011 to December 2020 comprises a high-statistics dataset of observations with an annual average of about 50 thunderstorm events [46].
@@ 5 Dependence on hadronic interaction generators
Since G3MT is a continuous sky monitor that monitors a huge volume of atmosphere, the dependence on hadronic interaction generators can be studied with many more such events. The G3MT has recorded a total of 487 significant thunderstorm events from April 2011 to December 2020 [46]. Among these, a list of eight major thunderstorm events, including the 1 December 2024, was prepared based on the maximum change of muon intensity and a relatively simple profile during the event. These events were analyzed following the same procedure as discussed in Section 4, except for the part that involved tracking of cloud movement, which would have required a less turbulent electric field for successful analysis. Table 2 lists the summary of these eight events with the total number of affected directions, the maximum change in muon intensity, the event duration, and the estimated thundercloud potentials derived for all nine combinations of hadronic interaction generators through Monte Carlo simulations in the subsequent columns. For each event, the relative difference in the thundercloud potential with respect to the SIBYLL-FLUKA combination is quoted in parentheses. To obtain a quantitative estimate of this dependence, a summary of the relative variations in thundercloud potential between these events for each combination is shown in Table 3. It is clear that the low-energy hadronic interaction generators display larger variations (14%), when compared to the high-energy generators (8%). Also, it should be noted that QGSJETII and GHEISHA provide systematically higher thundercloud potentials. Similarly, SIBYLL-FLUKA provide systematically lower thundercloud potentials.
@@ 6 Discussion
The G3MT is a ground-based instrument that is well suited for the study of thunderstorms. It is capable of measuring the variation in angular muon intensity (E1 GeV) with a precision of 0.1% on timescale of minutes. The G3MT had estimated an electric potential of 1.3 GV in one of the biggest thunderstorm events recorded on 1 December 2014, which confirmed an almost a century-old prediction by C.T.R. Wilson, superseding the earlier measurements by an order of magnitude carried out by sounding techniques. This estimation also offers an understanding the origin of high-energy photons detected in the terrestrial gamma-ray flashes by space probes. As discussed before, the estimations were based on Monte Carlo simulations, which are inherently dependent on hadronic interaction generators used in the development of EAS. It is shown that these generators produce minor differences in the magnitude of estimated thundercloud potential, even after tuning them with LHC data [47]. As mentioned before, these generators are built on assumptions based on a simple parton models associated with the Gribov-Regge multiple scattering. However, the implementation of physics treatment among these generators via different approaches leads to variation in the selection effects of pion, kaon, and muon production. As a result, notable differences are found among the predictions of these generators. Significant differences in the estimated thundercloud potential are observed among the nine combinations of hadronic interaction generators, derived from three generators each for the high- and low-energy segments, respectively. In the present work, eight major thunderstorm events observed by G3MT from April 2011 to December 2020 have been analyzed and their findings are reported. The high-energy generators cause a variation of 8%, whereas the low-energy generators produce variation of 14%. As the interactions below 80 GeV are handled by low-energy generators and a large majority of CRs are below this energy in the selected energy range of 10 GeV–10 TeV, the difference in low-energy generators dominate. The largest difference of 17% occurs for QGSJETII-GHEISHA with respect to the SIBYLL-FLUKA combination. Thus, the previously reported thundercloud potential of 1.3 GV is the most conservative estimate for the event of 1 December 2014 by using SIBYLL-FLUKA. Furthermore, this choice of SIBYLL-FLUKA generators also provides the lowest estimate of thundercloud potential for the other seven major thunderstorm events. This work provides a framework for estimating the thundercloud properties with the uncertainties introduced by different hadronic interaction generators used in CORSIKA, which are especially important for future such studies.
@@ Acknowledgement
We thank D.B. Arjunan, V. Jeyakumar, S. Kingston, K. Manjunath, S. Murugapandian, S. Pandurangan, B. Rajesh, V. Santhoshkumar, M.S. Shareef, C. Shobana, and R. Sureshkumar for their efforts in maintaining the GRAPES-3 experiment. The authors also express sincere thanks to the remaining members of the collaboration for their review and comments. We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI4002.

### v2

@@ Abstract
A potential difference of 1.3 Giga-Volts (GV) was inferred across a thundercloud using data from the GRAPES-3 muon telescope (G3MT) [1]. This was the first-ever estimation of gigavolt potential in thunderstorms, confirming prediction of C.T.R. Wilson almost a century ago. To infer the thundercloud potential required acceleration of muons in atmospheric electric field to be incorporated in the Monte Carlo simulation software CORSIKA. The G3MT records over 4 billion muons daily that are grouped into 169 directions covering 2.3 sr sky. This enabled changes as small as 0.1% in the muon intensity on minute timescale, caused by thunderstorms to be accurately measured. But that requires high statistics simulation of muons in thunderstorm electric fields. The CORSIKA offers a choice of several generators for low- (FLUKA, GHEISHA, and UrQMD) and high-energy (SIBYLL, EPOS-LHC, and QGSJETII) hadronic interactions. Since it is unclear which combination of the low- and high-energy generators provides the correct description of hadronic interactions, all nine combinations of generators were explored, and they yielded thundercloud potentials ranging from 1.3 GV to 1.6 GV for the event recorded on 1 December 2014. The result of SIBYLL-FLUKA combination yielded the lowest thundercloud potential of 1.3 GV was reported. Furthermore, another seven major thunderstorm events recorded between April 2011 and December 2020 were analyzed to measure the dependence of their thundercloud potential on the hadronic interaction generators. It is observed that the low-energy generators produce larger variation (14%) in thundercloud potential than the high-energy generators (8%). This probably reflects the fact that the GeV muons are predominantly produced in low-energy (80 GeV) interactions, which effectively magnifies the differences in the meson production cross-sections among the low-energy generators.
@@ 4 Thunderstorm events recorded by G3MT
The GeV muons detected by the G3MT, being relativistic particles, and combined with the fact that they almost traverse through the entire depth of the atmosphere, make them ideal candidates for thunderstorm studies. This may also explain the lack of success of sounding experiments to measure thundercloud potentials beyond 0.13 GV due to their limited coverage area and live-time [13]. In contrast, G3MT operates throughout the year with a sky coverage of 2.3 sr, which monitors a huge volume of the atmosphere with nearly 100% duty cycle. Although the G3MT has been operational for more than two decades, thunderstorm events could be studied in great detail after the installation of EFMs in April 2011. A collection of 487 significant thunderstorm events recorded from 2011 to December 2020 comprises a high-statistics dataset of observations with an annual average of about 50 thunderstorm events [46]. In this study, we analyze seven additional major thunderstorm events identified from the extended dataset corresponds to the above-mentioned period and estimate the corresponding thundercloud potentials. Several events exhibit gigavolt-scale potentials, indicating that such high-magnitude potentials are a characteristic feature of these thunderstorms. Since these estimates are derived through Monte Carlo simulations, it is crucial to quantitatively assess the differences among various hadronic event generators. To achieve this, a statistical analysis of the eight thunderstorm events is conducted to evaluate the overall variations introduced by different hadronic interaction models.
@@ 5 Dependence on hadronic interaction generators
From the extended database of 487 significant thunderstorm events from April 2011 to December 2020 [46], a list of eight major thunderstorm events, including the 1 December 2024, was prepared based on the maximum change of muon intensity and a relatively simple profile during the event. These events were analyzed following the same procedure as discussed in Section 4, except for the part that involved tracking of cloud movement, which would have required a less turbulent electric field for successful analysis. Table 2 lists the summary of these eight events with the total number of affected directions, the maximum change in muon intensity, the event duration, and the estimated thundercloud potentials derived for all nine combinations of hadronic interaction generators through Monte Carlo simulations in the subsequent columns. For each event, the relative difference in the thundercloud potential with respect to the SIBYLL-FLUKA combination is quoted in parentheses. To obtain a quantitative estimate of this dependence, a summary of the relative variations in thundercloud potential between these events for each combination is shown in Table 3. It is clear that the low-energy hadronic interaction generators display larger variations (14%), when compared to the high-energy generators (8%). Also, it should be noted that QGSJETII and GHEISHA provide systematically higher thundercloud potentials. Similarly, SIBYLL-FLUKA provide systematically lower thundercloud potentials.
The conclusion that the SIBYLL-FLUKA combination provides the lowest estimates of thundercloud potentials was initially derived from a comprehensive simulation dataset generated using CORSIKA v74001 with all nine available hadronic event generator combinations. However, as hadronic interaction models continue to be updated and refined, it is essential to assess the impact of these changes on the estimated thundercloud potentials. To evaluate the variability of the SIBYLL-FLUKA combination, we conducted additional simulations using the latest CORSIKA release, v78000, interfaced with SIBYLL 2.3e and FLUKA 2024.1. These simulations covered a thundercloud potential range of –3 GV to +3 GV, employing a coarser step size (0.25 GV instead of 0.1 GV in previous simulations) and reduced statistics (10% of the original dataset for each direction and thundercloud potential step). A comparative analysis of eight thunderstorm events revealed that the estimated thundercloud potentials differed by –(0.34.8)% between the two CORSIKA versions. The small mean variation (–0.3%) and standard deviation (4.8%) suggest that these differences lie within the measurement uncertainties (i.e., the 10% error on the estimated thundercloud potential, as shown in Figure 5 of [1]) and are statistically insignificant. Therefore, the thundercloud potentials derived using the SIBYLL-FLUKA combination, which consistently yielded the lowest values for the eight events analyzed in this study, remain robust and well-justified.
@@ 6 Discussion
The G3MT is a ground-based instrument that is well suited for the study of thunderstorms. It is capable of measuring the variation in angular muon intensity (E1 GeV) with a precision of 0.1% on timescale of minutes. The G3MT had estimated an electric potential of 1.3 GV in one of the biggest thunderstorm events recorded on 1 December 2014, which confirmed an almost a century-old prediction by C.T.R. Wilson, superseding the earlier measurements by an order of magnitude carried out by sounding techniques. This estimation also offers an understanding the origin of high-energy photons detected in the terrestrial gamma-ray flashes by space probes. As discussed before, the estimations were based on Monte Carlo simulations, which are inherently dependent on hadronic interaction generators used in the development of EAS. It is shown that these generators produce minor differences in the magnitude of estimated thundercloud potential, even after tuning them with LHC data [47]. As mentioned before, these generators are built on assumptions based on a simple parton models associated with the Gribov-Regge multiple scattering. However, the implementation of physics treatment among these generators via different approaches leads to variation in the selection effects of pion, kaon, and muon production. As a result, notable differences are found among the predictions of these generators. Significant differences in the estimated thundercloud potential are observed among the nine combinations of hadronic interaction generators, derived from three generators each for the high- and low-energy segments, respectively. In the present work, eight major thunderstorm events observed by G3MT from April 2011 to December 2020 have been analyzed and their findings are reported. The high-energy generators cause a variation of 8%, whereas the low-energy generators produce variation of 14%. As the interactions below 80 GeV are handled by low-energy generators and a large majority of CRs are below this energy in the selected energy range of 10 GeV–10 TeV, the difference in low-energy generators dominate. The largest difference of 17% occurs for QGSJETII-GHEISHA with respect to the SIBYLL-FLUKA combination. Given ongoing updates to hadronic event generators, we reassessed this conclusion using CORSIKA v78000 with SIBYLL 2.3e and FLUKA 2024.1. A comparison of eight thunderstorm events showed a deviation of –(0.34.8)%, which falls within the 10% measurement uncertainty (Figure 5 of [1]). These findings confirm that the SIBYLL-FLUKA combination remains the most consistent and conservative estimator of thundercloud potentials. Thus, the previously reported thundercloud potential of 1.3 GV is the most conservative estimate for the event of 1 December 2014 by using SIBYLL-FLUKA. Furthermore, this choice of SIBYLL-FLUKA generators also provides the lowest estimate of thundercloud potential for the other seven major thunderstorm events. This work provides a framework for estimating the thundercloud properties with the uncertainties introduced by different hadronic interaction generators used in CORSIKA, which are especially important for future such studies.
@@ 7 Summary
8 Future directions
The G3MT continues to record approximately 50 thunderstorm events annually. A comprehensive analysis of the entire dataset collected since its inception is essential for establishing the statistical distribution of thundercloud properties. This long-term study can offer insights into seasonal and interannual variations in thundercloud characteristics, contributing to a broader understanding of atmospheric electricity. Monte Carlo simulations remain indispensable for probing these properties using G3MT. This study has demonstrated the influence of hadronic event generators on the estimation of thundercloud properties. Future work may also explore the role of thundercloud geometry in simulations to refine these estimations further. Additionally, G3MT has been collecting high-resolution electric field measurements using four widely spaced EFMs over the past decade. This extensive dataset, combined with muon observations, presents a valuable opportunity for interdisciplinary research. Leveraging modern computational techniques, such as machine learning, to analyze the low-level data could provide deeper insights into the underlying physical processes, and develop predictive models for thunderstorm-related atmospheric phenomena.
@@ Acknowledgement
We thank D.B. Arjunan, V. Jeyakumar, S. Kingston, K. Manjunath, S. Murugapandian, S. Pandurangan, B. Rajesh, V. Santhoshkumar, M.S. Shareef, C. Shobana, and R. Sureshkumar for their efforts in maintaining the GRAPES-3 experiment. The authors also express sincere thanks to the remaining members of the collaboration for their review and comments. We acknowledge the support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI4002. We sincerely appreciate the referee’s thorough review and thoughtful suggestions, which have greatly contributed to the improvement of this manuscript. Their insightful feedback has helped refine our analysis and enhance the clarity of our presentation.